Please Explain in detail, draw illustrations to show this concept, legible please

An 850 kg undersea research station is a cube of edge length 5.00m. The station is tethered to the ocean floor by a 16.0m long cable in a location 30.0m deep. Given that sea water has a density of 1025 kg/m3, find the tension of the cable holding the station to the floor
Answer #1

here,

mass of cube , m = 850 kg

edge length , a = 5 m

volume , V = a^3 = 125 m^3

density , p = 1025 kg/m^3

the tension of the cable holding the station , T = Buoyant force - weight

T = p * V * g - m * g

T = 1025 * 125 * 9.81 - 850 * 9.81 N

T = 1.25 * 10^6 N

the tension of the cable holding the station is 1.25 * 10^6 N